# Unveiling Multiple Descents in Unsupervised Autoencoders

### v2

@@ Appendix B SNR Calculations
Let ${\boldsymbol{X}}\in\mathbb{R}^{D\times N}$ be the training data, with $\operatorname{rank}({\boldsymbol{X}})\geq m$. The reconstruction is given by
$\hat{{\boldsymbol{X}}}={\boldsymbol{W}}_{L}{\boldsymbol{W}}_{L-1}\cdots{\boldsymbol{W}}_{2}{\boldsymbol{W}}_{1}{\boldsymbol{X}}.$
@@
$\mathcal{L}=\|{\boldsymbol{X}}-\hat{{\boldsymbol{X}}}\|_{F}^{2}=\left\|{\boldsymbol{X}}-{\boldsymbol{W}}_{L}{\boldsymbol{W}}_{L-1}\cdots{\boldsymbol{W}}_{2}{\boldsymbol{W}}_{1}{\boldsymbol{X}}\right\|_{F}^{2}.$
If $\operatorname{rank}({\boldsymbol{W}}_{L-1}\cdots{\boldsymbol{W}}_{2})=d<m$, then the rank of $\hat{{\boldsymbol{X}}}$ is at most $d$. This means that ${\boldsymbol{W}}_{L-1}\cdots{\boldsymbol{W}}_{2}{\boldsymbol{W}}_{1}$ maps ${\boldsymbol{X}}$ into a subspace of dimension $d<m$. However, the optimal low-rank approximation of ${\boldsymbol{X}}$ that minimizes the reconstruction loss is achieved by projecting ${\boldsymbol{X}}$ onto the subspace spanned by its top $m$ singular vectors (the rank-$m$ approximation). By the Eckart-Young-Mirsky theorem (Eckart & Young, 1936),
$\mathcal{L}_{\text{optimal}}=\|{\boldsymbol{X}}-{\boldsymbol{X}}_{m}\|_{F}^{2},$
where ${\boldsymbol{X}}_{m}$ is the rank-$m$ approximation of ${\boldsymbol{X}}$. If $\operatorname{rank}({\boldsymbol{W}}_{L-1}\cdots{\boldsymbol{W}}_{2})<m$, the projection would not span the top $m$ singular vectors, leading to
@@ Appendix D Train Loss Results
Figure 12: Train loss of source data for the subspace data model.
@@ F.3 Double Descent Results For CNNs Trained on MNIST and CIFAR-10
In this section, we demonstrate that the double descent phenomenon can be reproduced in other unsupervised AE architectures. We employed the MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky et al., 2009) datasets and trained undercomplete CNNs as detailed in Figure LABEL:fig:cnn_ae_architecture. For the case of sample noise, the noise is added to $p\cdot 100\%$ of the images, and we present results for both MNIST and CIFAR-10, while all the other cases are presented using the MNIST dataset. For the feature noise scenario, noise is introduced to $p\cdot 100\%$ of the pixels of each image. To demonstrate the phenomenon with the presence of domain shift, the models are trained on the MNIST-M and MNIST datasets and tested on MNIST and MNIST-M, respectively. Results for model-wise double descent for varying amounts of sample and feature noise cases for the MNIST dataset are presented in Figure LABEL:fig:model_wise_sample_feature_noise_dd_mnist. Varying levels of sample noise and SNRs for the CIFAR-10 dataset are presented in Figure LABEL:fig:model_wise_sample_noise_dd_cifar.
In Figure LABEL:fig:denoising, we show the test and train loss results (top two sub-figures) for three different models (3, 5, and 60 channels) trained on MNIST with 50% sample noise and an SNR of -15 dB and find out that over-parameterized models can reduce the noise levels in an image. The smallest model, with 3 channels, is under-parameterized. The second model, within the critical regime, with 5 channels, performs poorly, while the third is over-parameterized, containing 60 channels. Interestingly, We noticed that even though our AE was not trained to remove noise (as in denoising AEs (Vincent et al., 2008, 2010)), over-parameterized models were able to reduce noise to some extent. In contrast, models within the critical regime performed significantly worse.
@@
$PSNR=10\cdot\log\left(\dfrac{R^{2}}{MSE({\boldsymbol{x}},f({\boldsymbol{x}}+\boldsymbol{\epsilon}))}\right),$
where ${\boldsymbol{x}}+\boldsymbol{\epsilon}$ represents the noisy image ($\boldsymbol{\epsilon}$ is the noise), and ${\boldsymbol{x}}$ is the clean version. This metric, expressed in decibels, allows us to evaluate how well each model cleans the images. As shown, the over-parameterized model consistently achieves the highest PSNR values (highlighted in bold green), while the poorly interpolating model, which primarily memorized noise, produces the lowest PSNR values (in red). In conclusion, over-parameterized models are capable of reducing noise when trained on noisy data, even without being explicitly tasked to do so.
We proceed by illustrating the impact of SNR on the test loss curve for both sample and feature noise scenarios in Figure LABEL:fig:model_wise_sample_feature_noise_snr_dd_mnist. As expected, low SNR values unveil double descent and increase test loss. We then investigate the effect of domain shifts between the training and testing datasets in two cases. First, models are trained on the MNIST dataset and tested on the MNIST-M dataset, as shown in Figure LABEL:fig:model_wise_domain_shift_dd_mnist. Second, models are trained on MNIST-M and tested on MNIST, as seen in Figure LABEL:fig:model_wise_domain_shift_dd_mnistm. In both cases, the model-wise double descent curve is observed.
@@ F.4 Double Descent Results for the Nonlinear Subspace Data Model
As in the subspace data model discussed in Subsection 3.1, we sample $N$ latent vectors $\{{\boldsymbol{z}}_{i}\}_{i=1}^{N}$ from a normal distribution and project them to a higher dimension using a random matrix ${\boldsymbol{H}}_{1}$. The key difference is the inclusion of nonlinear components ${\boldsymbol{z}}_{i}^{2}$ and ${\boldsymbol{z}}_{i}^{3}$, each projected to a higher dimensional space with different random matrices ${\boldsymbol{H}}_{2}$ and ${\boldsymbol{H}}_{3}$. To create contaminated setups of sample and feature noise, noise is added to $p\cdot 100\%$ of the data where $\beta$ controls the SNR:
$x_{i}=\begin{cases}\beta({\boldsymbol{H}}_{1}{\boldsymbol{z}}_{i}+{\boldsymbol{H}}_{2}{\boldsymbol{z}}_{i}^{2}+{\boldsymbol{H}}_{3}{\boldsymbol{z}}_{i}^{3})+\boldsymbol{\epsilon}_{i},&\text{with probability }p,\\ \beta({\boldsymbol{H}}_{1}{\boldsymbol{z}}_{i}+{\boldsymbol{H}}_{2}{\boldsymbol{z}}_{i}^{2}+{\boldsymbol{H}}_{3}{\boldsymbol{z}}_{i}^{3}),&\text{with probability }1-p.\end{cases}$
For the domain shift scenario, we divide the latent vectors into training and testing sets and use the same parameter ’$s$’ as described in Subsection 3.1 to control the shift between the train and test sets in the following manner: ${\boldsymbol{H}}_{j}^{{}^{\prime\prime}}={\boldsymbol{H}}_{j}+s\cdot{\boldsymbol{H}}^{{}^{\prime}}$ for $1\leq j\leq 3$ and get:
${\boldsymbol{x}}_{i}=\begin{cases}{\boldsymbol{H}}_{1}{\boldsymbol{z}}_{train}^{i}+{\boldsymbol{H}}_{2}({\boldsymbol{z}}_{train}^{i})^{2}+{\boldsymbol{H}}_{3}({\boldsymbol{z}}_{train}^{i})^{3},&\text{if }train,\\ {\boldsymbol{H}}_{1}^{{}^{\prime\prime}}{\boldsymbol{z}}_{test}^{i}+{\boldsymbol{H}}_{2}^{{}^{\prime\prime}}({\boldsymbol{z}}_{test}^{i})^{2}+{\boldsymbol{H}}_{3}^{{}^{\prime\prime}}({\boldsymbol{z}}_{test}^{i})^{3},&\text{if }test.\end{cases}$
For anomaly detection, clean samples are represented by $\beta({\boldsymbol{H}}_{1}{\boldsymbol{z}}_{i}+{\boldsymbol{H}}_{2}{\boldsymbol{z}}_{i}^{2}+{\boldsymbol{H}}_{3}{\boldsymbol{z}}_{i}^{3})$, with $p\cdot 100\%$ of them replaced by anomalies sampled from a normal distribution, as detailed in Subsection 3.1.

### v3

@@ Appendix B SNR Calculations
Let ${\bm{X}}\in\mathbb{R}^{D\times N}$ be the training data, with $\operatorname{rank}({\bm{X}})\geq m$. The reconstruction is given by
$\hat{{\bm{X}}}={\bm{W}}_{L}{\bm{W}}_{L-1}\cdots{\bm{W}}_{2}{\bm{W}}_{1}{\bm{X}}.$
@@
$\mathcal{L}=\|{\bm{X}}-\hat{{\bm{X}}}\|_{F}^{2}=\left\|{\bm{X}}-{\bm{W}}_{L}{\bm{W}}_{L-1}\cdots{\bm{W}}_{2}{\bm{W}}_{1}{\bm{X}}\right\|_{F}^{2}.$
If $\operatorname{rank}({\bm{W}}_{L-1}\cdots{\bm{W}}_{2})=d<m$, then the rank of $\hat{{\bm{X}}}$ is at most $d$. This means that ${\bm{W}}_{L-1}\cdots{\bm{W}}_{2}{\bm{W}}_{1}$ maps $\bm{X}$} into a subspace of dimension $d<m$. However, the optimal low-rank approximation of $\bm{X}$} that minimizes the reconstruction loss is achieved by projecting $\bm{X}$} onto the subspace spanned by its top $m$ singular vectors (the rank-$m$ approximation). By the Eckart-Young-Mirsky theorem (Eckart & Young, 1936),
$\mathcal{L}_{\text{optimal}}=\|{\bm{X}}-{\bm{X}}_{m}\|_{F}^{2},$
where ${\bm{X}}_{m}$ is the rank-$m$ approximation of $\bm{X}$}. If $\operatorname{rank}({\bm{W}}_{L-1}\cdots{\bm{W}}_{2})<m$, the projection would not span the top $m$ singular vectors, leading to
@@
Remark C.3 (Generalization for noisy test data).
Throughout the paper, we trained non-linear models on contaminated datasets, drawn from $\mathcal{P}_{{\bm{x}}}$ and tested on clean data sampled from $\mathcal{P}_{\tilde{{\bm{x}}}}$ to isolate the effect on the learned model, showing noise memorization (high test loss) versus signal learning (low test loss) and observed model, epoch, and sample-wise multiple descents. We followed a similar approach as in supervised learning, where models are trained with noisy labels and tested on clean ones (in our unsupervised scenario, we trained on noisy data and tested on clean data, similar to (Lupidi et al., 2023)). We also showed that for the case of linear models, the test loss decreases as model size increases. If we define the generalization loss differently than the definition in the paper, to account for noise in the test data, then the test loss does not have to monotonically decrease when $m$ increases.
[…]
$=\mathbb{E}\Big{[}\big{\|}\mathbf{U}_{m}\mathbf{U}_{m}^{\top}(\tilde{{\bm{x}}}+{\bm{n}})-\tilde{{\bm{x}}}\big{\|}_{2}^{2}\Big{]}$
$=\mathbb{E}\Big{[}\big{\|}(\mathbf{U}_{m}\mathbf{U}_{m}^{\top}-\mathbb{I})\tilde{{\bm{x}}}+\mathbf{U}_{m}\mathbf{U}_{m}^{\top}{\bm{n}}\big{\|}_{2}^{2}\Big{]}$
$=\mathbb{E}\Big{[}\big{\|}(\mathbb{I}-\mathbf{U}_{m}\mathbf{U}_{m}^{\top})\tilde{{\bm{x}}}\big{\|}_{2}^{2}+\big{\|}\mathbf{U}_{m}\mathbf{U}_{m}^{\top}{\bm{n}}\big{\|}_{2}^{2}+2\big{\langle}(\mathbb{I}-\mathbf{U}_{m}\mathbf{U}_{m}^{\top})\tilde{{\bm{x}}},\mathbf{U}_{m}\mathbf{U}_{m}^{\top}{\bm{n}}\big{\rangle}\Big{]}$
$=\mathbb{E}\Big{[}\big{\|}(\mathbb{I}-\mathbf{U}_{m}\mathbf{U}_{m}^{\top})\tilde{{\bm{x}}}\big{\|}_{2}^{2}\Big{]}+\mathbb{E}\Big{[}\big{\|}\mathbf{U}_{m}\mathbf{U}_{m}^{\top}{\bm{n}}\big{\|}_{2}^{2}\Big{]}$
$=\mathrm{Tr}\Big{[}(\mathbb{I}-\mathbf{U}_{m}\mathbf{U}_{m}^{\top})\Sigma_{\tilde{{\bm{x}}}}\Big{]}+\mathrm{Tr}\Big{[}\mathbf{U}_{m}\mathbf{U}_{m}^{\top}\Sigma_{\bm{n}}\Big{]}$
$=\underbrace{\mathrm{Tr}\Big{[}(\mathbb{I}-\mathbf{U}_{m}\mathbf{U}_{m}^{\top})\Sigma_{\tilde{{\bm{x}}}}\Big{]}}_{\text{clean error}}+\underbrace{\mathrm{Tr}\Big{[}\mathbf{U}_{m}\mathbf{U}_{m}^{\top}\Sigma_{\bm{n}}\Big{]}}_{\text{projected noise contribution}}.$
We used the assumption that the expectation of the noise is zero. When $m$ increases, the first term decreases while the second term increases, implying that increasing $m$ does not have to monotonically decrease this definition of the generalization error.
@@ F.3 Double Descent Results For CNNs Trained on MNIST and CIFAR-10
In this section, we demonstrate that the double descent phenomenon can be reproduced in other unsupervised AE architectures. We employed the MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky et al., 2009) datasets and trained undercomplete CNNs as detailed in Figure LABEL:fig:cnn_ae_architecture. For the case of sample noise, the noise is added to $p\cdot 100\%$ of the images, and we present results for both MNIST and CIFAR-10, while all the other cases are presented using the MNIST dataset. For the feature noise scenario, noise is introduced to $p\cdot 100\%$ of the pixels of each image. To demonstrate the phenomenon with the presence of domain shift, the models are trained on the MNIST-M and MNIST datasets and tested on MNIST and MNIST-M, respectively. Results for model-wise double descent for varying amounts of sample and feature noise cases for the MNIST dataset are presented in Figures LABEL:fig:model_wise_sample_noise_mnist, LABEL:fig:model_wise_sample_feature_noise_dd_mnist respectively. Varying levels of sample noise and SNRs for the CIFAR-10 dataset are presented in Figure LABEL:fig:model_wise_sample_noise_dd_cifar.
In Figure LABEL:fig:denoising, we show the test and train loss results (top two sub-figures) for three different models (3, 5, and 60 channels) trained on MNIST with 50% sample noise and an SNR of -15 dB and find out that over-parameterized models can reduce the noise levels in an image. The smallest model, with 3 channels, is under-parameterized. The second model, within the critical regime, with 5 channels, performs poorly, while the third is over-parameterized, containing 60 channels. Interestingly, We noticed that even though our AE was not trained to remove noise (as in denoising AEs (Vincent et al., 2008; 2010)), over-parameterized models were able to reduce noise to some extent. In contrast, models within the critical regime performed significantly worse.
@@
$PSNR=10\cdot\log\left(\dfrac{R^{2}}{MSE({\bm{x}},f({\bm{x}}+\bm{\epsilon}))}\right),$
where ${\bm{x}}+\bm{\epsilon}$ represents the noisy image ($\bm{\epsilon}$ is the noise), and $\bm{x}$} is the clean version. This metric, expressed in decibels, allows us to evaluate how well each model cleans the images. As shown, the over-parameterized model consistently achieves the highest PSNR values (highlighted in bold green), while the poorly interpolating model, which primarily memorized noise, produces the lowest PSNR values (in red). In conclusion, over-parameterized models are capable of reducing noise when trained on noisy data, even without being explicitly tasked to do so.
We proceed by illustrating the impact of SNR on the test loss curve for both sample and feature noise scenarios in Figures LABEL:fig:model_wise_sample_noise_snr_mnist, LABEL:fig:model_wise_sample_feature_noise_snr_dd_mnist respectively. As expected, low SNR values unveil double descent and increase test loss. We then investigate the effect of domain shifts between the training and testing datasets in two cases. First, models are trained on the MNIST dataset and tested on the MNIST-M dataset, as shown in Figure LABEL:fig:model_wise_domain_shift_mnist_mnistm. Second, models are trained on MNIST-M and tested on MNIST, as seen in Figure LABEL:fig:model_wise_domain_shift_dd_mnistm. In both cases, the model-wise double descent curve is observed.
@@ F.4 Double Descent Results for the Nonlinear Subspace Data Model
As in the subspace data model discussed in Subsection 3.1, we sample $N$ latent vectors $\{{\bm{z}}_{i}\}_{i=1}^{N}$ from a normal distribution and project them to a higher dimension using a random matrix ${\bm{H}}_{1}$. The key difference is the inclusion of nonlinear components ${\bm{z}}_{i}^{2}$ and ${\bm{z}}_{i}^{3}$, each projected to a higher dimensional space with different random matrices ${\bm{H}}_{2}$ and ${\bm{H}}_{3}$. To create contaminated setups of sample and feature noise, noise is added to $p\cdot 100\%$ of the data where $\beta$ controls the SNR:
$x_{i}=\begin{cases}\beta({\bm{H}}_{1}{\bm{z}}_{i}+{\bm{H}}_{2}{\bm{z}}_{i}^{2}+{\bm{H}}_{3}{\bm{z}}_{i}^{3})+\bm{\epsilon}_{i},&\text{with probability }p,\\ \beta({\bm{H}}_{1}{\bm{z}}_{i}+{\bm{H}}_{2}{\bm{z}}_{i}^{2}+{\bm{H}}_{3}{\bm{z}}_{i}^{3}),&\text{with probability }1-p.\end{cases}$
For the domain shift scenario, we divide the latent vectors into training and testing sets and use the same parameter ‘$s$’ as described in Subsection 3.1 to control the shift between the train and test sets in the following manner: ${\bm{H}}_{j}^{{}^{\prime\prime}}={\bm{H}}_{j}+s\cdot{\bm{H}}^{{}^{\prime}}$ for $1\leq j\leq 3$ and get:
${\bm{x}}_{i}=\begin{cases}{\bm{H}}_{1}{\bm{z}}_{train}^{i}+{\bm{H}}_{2}({\bm{z}}_{train}^{i})^{2}+{\bm{H}}_{3}({\bm{z}}_{train}^{i})^{3},&\text{if }train,\\ {\bm{H}}_{1}^{{}^{\prime\prime}}{\bm{z}}_{test}^{i}+{\bm{H}}_{2}^{{}^{\prime\prime}}({\bm{z}}_{test}^{i})^{2}+{\bm{H}}_{3}^{{}^{\prime\prime}}({\bm{z}}_{test}^{i})^{3},&\text{if }test.\end{cases}$
For anomaly detection, clean samples are represented by $\beta({\bm{H}}_{1}{\bm{z}}_{i}+{\bm{H}}_{2}{\bm{z}}_{i}^{2}+{\bm{H}}_{3}{\bm{z}}_{i}^{3})$, with $p\cdot 100\%$ of them replaced by anomalies sampled from a normal distribution, as detailed in Subsection 3.1.